# Parameter estimation in fluid flow models from undersampled frequency space data

### v1

@@ 2.2.3 Objective function
$\bm{\hat{\theta}}=\operatorname*{\operatorname{argmin}}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2\sigma_{v}^{2}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}\left(\left[\bm{V}^{n}-\mathcal{H}(\bm{u}_{\bm{\theta}}^{n})\right]_{s}\right)^{2}+\frac{1}{2}||\bm{\theta}-\bm{\theta}^{0}||_{(\bm{P}^{0})^{-1}}^{2}$ (8)
@@
where $\bm{V}^{k}\in\mathbb{R}^{N}$ are the measurements of the velocity provided by the reconstruction of the PC-MRI acquisition and $\mathcal{H}:[H_{1}(\Omega)]^{3}\to\mathbb{R}^{N}$ is the observation operator, which is applied to the result $\bm{u}_{\bm{\theta}}^{n}$ of the forward model for a set of model parameters $\bm{\theta}\in\mathbb{R}^{p}$ at the time corresponding to the measurement $n$. In this standard inverse problem for velocities, this observation operator usually corresponds to the interpolation from the model’s (usually finer) mesh to an array of measured velocities at each image voxel. In fact, [.$[.]_{s}$ denotes the $s$-th vector element (or voxel in the case of image data), therefore summing over all the vector elements. $\bm{\theta}^{0}$ is the initial guess for the parameters with its covariance matrix $\bm{P}^{0}$, both of which are given by the user but they are assumed to be known for the numerical inverse solver. $\sigma_{v}$ denotes the standard deviation of the noise on the velocity measurements, which we also assume to be (approximately) known.
@@
This inverse problem has been commonly applied to the estimation of boundary conditions in hemodynamics. For a thorough review, see [1].
@@
In [9], an alternative formulation of this objective function was proposed to account for aliasing artifacts present in velocity MRI data. The functional in this case is
$\bm{\hat{\theta}}$ $=$ $\operatorname*{\operatorname{argmin}}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2\sigma_{M}^{2}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}\Big{(}[\Re(\bm{M}_{meas}^{n})-|\bm{M}_{meas}^{n}|\cos(\phi_{back}+\mathcal{H}(\bm{u}_{\bm{\theta}}^{n})\frac{\pi}{venc})]_{s}\Big{)}^{2}$ (9)
$+\frac{1}{2\sigma_{M}^{2}}\Big{(}[\Im(\bm{M}_{meas}^{n})-|\bm{M}_{meas}^{n}|\sin(\phi_{back}+\mathcal{H}(\bm{u}^{n}_{\bm{\theta}})\frac{\pi}{venc}]_{s})\Big{)}^{2}+\frac{1}{2}||\bm{\theta}-\bm{\theta}^{0}||_{(\bm{P}^{0})^{-1}}^{2}$
@@
$=$ $\operatorname*{\operatorname{argmin}}_{\bm{\theta}\in\mathbb{R}^{p}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}\frac{|\bm{M}_{meas}^{n}|_{s}^{2}}{\sigma_{M}^{2}}\Big{(}1-\cos(\frac{\pi}{venc}([\bm{V}^{n}-\mathcal{H}(\bm{u}^{n}_{\bm{\theta}})))]_{s}\Big{)}+\frac{1}{2}||\bm{\theta}-\bm{\theta}^{0}||_{(\bm{P}^{0})^{-1}}^{2}$
where $\bm{M}^{n}_{meas}$ are measurements of the complex magnetization and $\sigma_{M}$ is the standard deviation of the noise for each of the (complex magnetization) real/imaginary measurements. This formulation assumes that the magnitude $|\bm{M}^{n}_{meas}|$ of the magnetization is known in order to formulate the problem in terms of the measured and observed velocities.
Still, the aforementioned cost functions do not account for artifacts originating from the frequency undersampling of the data. Therefore, as we will observe in the numerical examples, the error in the data and hence in the estimated parameters in the inverse problem drastically grows when increasing the undersampling. Here, we postulate to solve this issue by formulating the parameter estimation problem as it is done in CS through a data fidelity term in terms of the k-space, therefore leading to the minimization problem:
@@
$\bm{\hat{\theta}}=\operatorname*{\operatorname{argmin}}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2\sigma_{y}^{2}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}([\Re(\bm{Y}^{n}-\mathcal{H}_{\mathcal{F}}(\bm{u}_{\bm{\theta}}^{n})]_{s})^{2}+([\Im(\bm{Y}^{n}-\mathcal{H}_{\mathcal{F}}(\bm{u}_{\bm{\theta}}^{n}))]_{s})^{2}+\frac{1}{2}||\bm{\theta}-\bm{\theta}^{0}||_{(\bm{P}^{0})^{-1}}^{2}$ (10)
@@
$\mathcal{H_{\mathcal{F}}}(u)=\mathcal{F}\left(\bm{M}\odot e^{i\frac{\pi}{venc}\bm{u}+\phi_{back}}\right)\odot\bm{S}$ (11)
where $\bm{S}\in\mathbb{R}^{N}$ is the sampling mask, i.e. with the entries corresponding to the sampled voxels set to $1$, and the others to 00. The $\bm{Y}^{n}$ are measurements in frequency space and $\sigma_{y}$ is the standard deviation of the noise on the frequency-space measurements, which is assumed to be Gaussian and zero-mean. The use of this observation operator requires that the magnitude of the magnetization, or an approximation thereof, is known.
@@ 2.2.4 Parameter estimation with Kalman filtering
There are different approaches to solving the optimization problems described above. Adjoint-based variational data assimilation approaches fit the entire ensemble of measurements to the ensemble of the observations at the matching time steps. However, this poses significant storage requirements, as the entire trajectory needs to be stored. On the other hand, sequential data assimilation assimilates the information from each new measurement at the point that measurement occurs during one forward pass of the model. This leads to a sequential improvement of the estimation over the duration of the inverse problem as more measurements are considered.
We have chosen to use the Reduced Order Unscented Kalman Filter (ROUKF) in order to solve the optimization problem (10), as it is computationally tractable and has been successfully used in blood-flow problems and other time-resolved problems due to its recursivity.
@@
$\bm{\Gamma}^{n}=\bm{V}^{n}-\mathcal{H}(\bm{u}^{n})$ (12)
@@
$\bm{\Gamma}^{n}=\frac{1}{\sqrt{2}}|\bm{M}(t^{n})|\sin\left(\frac{\pi}{venc}\cdot\left(\bm{V}^{n}-\mathcal{H}(\bm{u}^{n})\right)\right)$
where the factor $\frac{\sqrt{2}venc}{\pi}|\bm{M}(t^{n})|^{-1}$ is added to ensure the equivalence of this innovation with Equation (12) in the case of high $venc$. For more details, see [9].
As done for the first cost function, assuming Gaussianity in complex magnetization measurements, the innovation for the k-space based objective function (10) is defined by:
$\bm{\Gamma}_{n}=\begin{bmatrix}\Re(\bm{Y}^{n})-\Re(\mathcal{H}_{\mathcal{F}}(\bm{u}^{n}))\\ \Im(\bm{Y}^{n})-\Im(\mathcal{H}_{\mathcal{F}}(\bm{u}^{n}))\end{bmatrix}.$ (13)
Moreover, we remark that we are adapting the Kalman filter as described in [10]. Instead of simplex sigma points (which involve $p+1$ solutions of the forward problem), we are using canonical sigma points, which involve $2p$ solutions of the forward problem but they have shown in our simulations better performance than their simplex counterparts. In particular, the results using both types of sigma points deviate more the larger the number of parameters to be estimated. Also, the results with simplex points change depending on the enumeration of the parameters during the estimation process, which is not the case with canonical points. Additionally, to ensure positivity of the parameters, we are reparameterizing such that $\bm{\theta}=\bm{\theta}^{0}2^{\nu}$ with $\bm{\theta}^{0}$ the initial guess for the parameters and the filter being applied to $\nu$, as is also done in [9].
@@ 3.1.1 Forward problem setup
We use the same model as is used in [9]. We consider a geometry of the lumen of the ascending and descending aorta including the outlets of the brachiocephalic artery, left common carotid artery, and left subclavian artery, as depicted in Figure 1. This geometry serves as the domain for the forward model.
@@
The boundary of the geometry consists of six different boundaries: $\Gamma_{in}$ being the inlet boundary in the ascending aorta, $\Gamma_{w}$ the arterial wall, and the remaining boundaries $\Gamma_{l}$ for $l=1,\cdots,4$ respresenting the outlets.
We model the blood flow in this domain with the incompressible Navier-Stokes equations for the velocity $\bm{u}(\bm{x},t)\in\mathbb{R}^{3}$ and the pressure $p(\bm{x},t)\in\mathbb{R}$:
$\begin{cases}\rho\frac{\partial\bm{u}}{\partial t}+\rho(\bm{u}\cdot\nabla)\bm{u}-\mu\Delta\bm{u}+\nabla p=0\text{ in }\Omega\\ \nabla\cdot\bm{u}=0\text{ in }\Omega\\ \bm{u}=\bm{u}_{in}\text{ on }\Gamma_{in}\\ \bm{u}=\bm{0}\text{ on }\Gamma_{w}\\ \mu\frac{\partial\bm{u}}{\partial\bm{n}}-p\bm{n}=-P_{l}(t)\bm{n}\text{ on }\Gamma_{l}\end{cases}$ (14)
@@
[…]
@@
The inflow $\bm{u}_{in}$ is defined as
$\bm{u}_{in}=-Uf(t)\bm{n}$
@@
The forward problem is solved using a semi-implicit 3D-0D coupling scheme as seen in [9]. The full algorithm is detailed in the appendix.
@@ 3.1.2 Synthetic measurements
The forward solution is generated with a time step of $dt=1ms$ and undersampled in time to $dt_{meas}=15ms$, leading to a total of 56 measurements. From the solution of the forward problem, we simulate a PC-MRI acquisition by subsampling into a rectangular measurement mesh with a resolution of $[2mm,2mm,2mm]$ and then applying the process described in Section 2.1 with a $venc$ of double the maximal velocity. The magnitude is modelled as
$M(\bm{x})=\begin{cases}1.0\text{ if }\bm{x}\text{ is in the lumen of the vessel}\\ 0.5\text{ otherwise.}\end{cases}$ (16)
Finally a complex Gaussian noise $\bm{\epsilon}\in\mathbb{C}^{N}$ is added with a signal-to-noise ratio (SNR) of 15. Fifty independent realizations of the noise were generated.
For comparison, we reconstructed velocity measurements from these synthetic measurements using the Berkeley Advanced Reconstruction Toolbox (BART)[11]. BART is a command-line-based software that provides a flexible framework of compressed sensing methods, as well as tools for simulation, pre-processing, and image reconstruction, providing a multitude of different regularization options. In this work, we have used this toolbox for compressed sensing reconstructions of the velocities, using total variation in time as for the regularization.
Next, the sampling mask is applied to these simulated frequency space measurements. We take a 2D subsampled mask in the $x-y$-plane and sample fully in the $z$-direction as in [4]. We consider different subsampling rates $R=\frac{N_{sampled}}{N_{total}}={8,16,32}$, with two different masks: the pseudo-spiral mask and the pseudo-random Gaussian mask, which is sampled according to a Gaussian probability distribution, as shown in Figure 2. For the pseudo-spiral mask, the points are placed evenly on a cartesian grid along a spiral with six turns and a final radius reaching the edge of the mask.
Additionally, we require measurements of the magnitude itself.
@@ 3.1.3 Inverse problem setup
We consider two different initial guesses for these parameters, as in [9]
@@
The initial standard deviation for the reparameterized parameters was set to $0.5$, i.e. $P^{0}=0.5\mathbb{I}$, meaning that the prior models that there is $\approx 95\%$ probability that the target value will lie between the range half/twice the initial guess. The standard deviation of the noise was estimated from the initial time step by computing the standard deviation of
$\bm{Y}^{0}-\bm{M}(t^{0})\odot\exp\left(i\phi_{back}(t^{0}))\right)\odot\bm{S}$
@@ 3.2.1 Flow phantom measurements
We used data from a flow experiment on a phantom of the carotid artery reported in [4]. The phantom is made of a distensible silicon and is suspended in water. It takes the shape of a bifurcating tube, to simulate a carotid artery, and a backflow tube. A pump generates a pulsatile flow with a rate of roughy 60 bpm, simulating a cardiac cycle lasting 1s.
The 4D Flow MRI scan was performed with a 3T Ingenia scanner by Philips Healthcare. All three velocity directions were acquired, plus one acquisition with no encoding gradient to acquire the background phase. The scan parameters were set to TR = 8.9ms, TE = 4.5ms, FA = 8, $venc$=150cm/s. The matrix size was [160, 160, 40] with a spatial resolution of [0.8mm, 0.8mm, 0.8mm] and a temporal resolution of 19 frames per cardiac cycle. The scan was accelerated with an acceleration factor of R=2. The scan used a 32-channel coil leading to 15 independent coil measurements, resulting in a total matrix size of $[160,160,40,4,15,19]$.

### v2

@@ 2.2.3 Objective function
[…]
@@
where $\boldsymbol{V}^{k}\in\mathbb{R}^{N}$ are the measurements of the velocity provided by the reconstruction of the PC-MRI acquisition and $\mathcal{H}:[H_{1}(\Omega)]^{3}\to\mathbb{R}^{N}$ is the observation operator, which is applied to the result $\boldsymbol{u}_{\boldsymbol{\theta}}^{n}$ of the forward model for a set of model parameters $\boldsymbol{\theta}\in\mathbb{R}^{p}$ at the time corresponding to the measurement $n$. In this standard inverse problem for velocities, this observation operator usually corresponds to the interpolation from the model’s (usually finer) mesh to an array of measured velocities at each image voxel. In fact, $_{s}$ denotes the $s$-th vector element (or voxel in the case of image data), therefore summing over all the vector elements. $\boldsymbol{\theta}^{0}$ is the initial guess for the parameters with its covariance matrix $\boldsymbol{P}^{0}$, both of which are given by the user but are assumed to be known for the numerical inverse solver. $\sigma_{v}$ denotes the standard deviation of the noise on the velocity measurements, which we also assume to be (approximately) known.
@@
This inverse problem has been commonly applied to the estimation of boundary conditions in hemodynamics. For a thorough review, see [nolte2022review].
@@
In [garay2022], an alternative formulation of this objective function was proposed to account for aliasing artifacts present in velocity MRI data. The functional in this case is
@@
$\boldsymbol{\hat{\theta}}$ $=$ $\text{argmin}_{\boldsymbol{\theta}\in\mathbb{R}^{p}}\frac{1}{2\sigma_{M}^{2}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}\Big(\Big([\Re(\boldsymbol{M}_{meas}^{n})$ (9)
$-|\boldsymbol{M}_{meas}^{n}|\cos(\phi_{back}+\mathcal{H}(\boldsymbol{u}_{\boldsymbol{\theta}}^{n})\frac{\pi}{venc})]_{s}\Big)^{2}$
$+\Big([\Im(\boldsymbol{M}_{meas}^{n})-|\boldsymbol{M}_{meas}^{n}|\sin(\phi_{back}+\mathcal{H}(\boldsymbol{u}^{n}_{\boldsymbol{\theta}})\frac{\pi}{venc}]_{s})\Big)^{2}\Big)$
[…]
@@
$=$ $\text{argmin}_{\boldsymbol{\theta}\in\mathbb{R}^{p}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}\frac{|\boldsymbol{M}_{meas}^{n}|_{s}^{2}}{\sigma_{M}^{2}}\Big(1-\cos(\frac{\pi}{venc}([\boldsymbol{V}^{n}-\mathcal{H}(\boldsymbol{u}^{n}_{\boldsymbol{\theta}})))]_{s}\Big)$
[…]
where $\boldsymbol{M}^{n}_{meas}$ are measurements of the complex magnetization and $\sigma_{M}$ is the standard deviation of the noise for each of the (complex magnetization) real/imaginary measurements. This formulation assumes that the magnitude $|\boldsymbol{M}^{n}_{meas}|$ of the magnetization is known in order to formulate the problem in terms of the measured and observed velocities.
Still, the aforementioned cost functions do not account for artifacts originating from the frequency undersampling of the data. Therefore, as we will observe in the numerical examples, the error in the data and hence in the estimated parameters in the inverse problem drastically grows when increasing the undersampling. Here, we propose to solve this issue by formulating the parameter estimation problem as it is done in CS through a data fidelity term in terms of the k-space, therefore leading to the minimization problem:
$\boldsymbol{\hat{\theta}}$ $=$ $\text{argmin}_{\boldsymbol{\theta}\in\mathbb{R}^{p}}\frac{1}{2\sigma_{y}^{2}}\sum_{n=1}^{N_{T}}\sum_{s=1}^{N}([\Re(\boldsymbol{Y}^{n}-\mathcal{H}_{\mathcal{F}}(\boldsymbol{u}_{\boldsymbol{\theta}}^{n})]_{s})^{2}+([\Im(\boldsymbol{Y}^{n}-\mathcal{H}_{\mathcal{F}}(\boldsymbol{u}_{\boldsymbol{\theta}}^{n}))]_{s})^{2}$ (10)
[…]
@@
$\mathcal{H_{\mathcal{F}}}(u)=\mathcal{F}\left(\boldsymbol{M}\odot e^{i\frac{\pi}{venc}\boldsymbol{u}+\phi_{back}}\right)\odot\boldsymbol{S}$ (11)
where $\boldsymbol{S}\in\mathbb{R}^{N}$ is the sampling mask, i.e. with the entries corresponding to the sampled voxels set to $1$, and the others to 0. The $\boldsymbol{Y}^{n}$ are measurements in frequency space and $\sigma_{y}$ is the standard deviation of the noise on the frequency-space measurements, which is assumed to be Gaussian and zero-mean. The use of this observation operator requires that the magnitude of the magnetization as well as the background phase, or an approximation thereof, is known.
@@ 2.2.4 Parameter estimation with Kalman filtering
There are various approaches to solving the optimization problems described above. Adjoint-based variational data assimilation approaches fit the entire ensemble of measurements to the ensemble of the observations at the matching time steps. However, this poses significant storage requirements, as the entire trajectory needs to be stored. On the other hand, sequential data assimilation assimilates the information from each new measurement at the point that measurement occurs during one forward pass of the model. This leads to a sequential improvement of the estimation over the duration of the inverse problem as more measurements are considered.
We have chosen to use the Reduced Order Unscented Kalman Filter (ROUKF)[moireau-chapelle-11] in order to solve the optimization problem (10), as it is computationally tractable and has been successfully used in blood-flow problems and other time-resolved problems due to its recursivity. ROUKF has been successfully employed for parameter estimation from (already reconstructed) MR images in this context before, alas not with the raw frequency space data [arthurs2020flexible, imperiale_sequential_2021].
@@
$\boldsymbol{\Gamma}^{n}=\boldsymbol{V}^{n}-\mathcal{H}(\boldsymbol{u}^{n})$ (12)
@@
$\boldsymbol{\Gamma}^{n}=\frac{1}{\sqrt{2}}|\boldsymbol{M}(t^{n})|\sin\left(\frac{\pi}{venc}\cdot\left(\boldsymbol{V}^{n}-\mathcal{H}(\boldsymbol{u}^{n})\right)\right)$
where the factor $\frac{\sqrt{2}venc}{\pi}|\boldsymbol{M}(t^{n})|^{-1}$ is added to ensure the equivalence of this innovation with Equation (12) in the case of high vencvenc. For more details, see [garay2022].
As done for the first cost function, assuming a Gaussian distribution in the noise of the complex magnetization measurements, the innovation for the k-space based objective function (10) is defined by:
$\boldsymbol{\Gamma}_{n}=\begin{bmatrix}\Re(\boldsymbol{Y}^{n})-\Re(\mathcal{H}_{\mathcal{F}}(\boldsymbol{u}^{n}))\\ \Im(\boldsymbol{Y}^{n})-\Im(\mathcal{H}_{\mathcal{F}}(\boldsymbol{u}^{n}))\end{bmatrix}$. (13)
Moreover, we remark that we are adapting the Kalman filter as described in [moireau-chapelle-11]. Instead of simplex sigma points (which involve $p+1$ solutions of the forward problem), we are using canonical sigma points, which involve $2p$ solutions of the forward problem but have shown in our simulations better performance than their simplex counterparts. In particular, the results using both types of sigma points deviate more the larger the number of parameters to be estimated. Also, the results with simplex points change depending on the enumeration of the parameters during the estimation process, which is not the case with canonical points.
Additionally, to ensure positivity of the parameters, we are reparameterizing such that $\boldsymbol{\theta}=\boldsymbol{\theta}^{0}2^{\nu}$ with $\boldsymbol{\theta}^{0}$ the initial guess for the parameters and the filter being applied to $\nu$, as is also done in [garay2022]. The initial value of $\nu$ is set to $\boldsymbol{0}$.
@@ 3.1.1 Forward problem setup
We use the same model as is used in [garay2022]. We consider a geometry of the lumen of the ascending and descending aorta including the outlets of the brachycephalic artery, left common carotid artery, and left subclavian artery, as depicted in Figure 1. This geometry serves as the domain for the forward model. The geometry was discretized with unstructured trapezoidal elements with a total of 20,916 points.
@@
The boundary of the geometry consists of six different boundaries: $\Gamma_{in}$ being the inlet boundary in the ascending aorta, $\Gamma_{w}$ the arterial wall, and the remaining boundaries $\Gamma_{l}$ for $l=1,\cdots,4$ representing the outlets.
We model the blood flow in this domain with the incompressible Navier-Stokes equations for the velocity $\boldsymbol{u}(\boldsymbol{x},t)\in\mathbb{R}^{3}$ and the pressure $p(\boldsymbol{x},t)\in\mathbb{R}$:
[…]
@@
$\begin{cases}P_{l}=R_{p,l}Q_{l}+\pi_{l}\\ Q_{l}=\int_{\Gamma_{l}}\boldsymbol{u}\cdot\boldsymbol{n}dx\\ C_{d,l}\frac{d\pi_{l}}{dt}+\frac{\pi_{l}}{R_{d,l}}=Q_{l}\end{cases}$ (15)
@@
The inflow $\boldsymbol{u}_{in}$ is defined as
$\boldsymbol{u}_{in}=-Uf(t)\boldsymbol{n}$
@@
The forward problem is solved using an in-house finite elements solver, with a semi-implicit 3D-0D coupling scheme as in [garay2022] and using $P1$ elements for both the velocity and the pressure. The full algorithm is detailed in the appendix.
@@ 3.1.2 Synthetic measurements
The forward solution is generated with a time step of $dt=1ms$ and undersampled in time to $dt_{meas}=15ms$, leading to a total of 56 measurements. From the solution of the forward problem, we simulate a PC-MRI acquisition by subsampling into a rectangular measurement mesh with a resolution of $[2mm,2mm,2mm]$ and then applying the process described in Section 2.1 with a vencvenc of double the maximal velocity. The magnitude is modelled as
[…]
and the background phase was set to an arbitrary constant value of $\phi_{back}=7.5\cdot 10^{-2}rad$. Finally a complex Gaussian noise $\boldsymbol{\epsilon}\in\mathbb{C}^{N}$ is added with a signal-to-noise ratio (SNR) of 15. Fifty independent realizations of the noise were generated.
For comparison, we reconstructed velocity measurements from these synthetic measurements using the Berkeley Advanced Reconstruction Toolbox (BART)[bart2015software]. BART is a command-line-based software that provides a flexible framework of compressed sensing methods, as well as tools for simulation, pre-processing, and image reconstruction, providing a multitude of different regularization options. In this work, we have used this toolbox for compressed sensing reconstructions of the velocities, using total variation in time as for the regularization. Examples of the reconstructed velocities, using different masks and acceleration factors, are shown in Figure 4.
Next, the sampling mask is applied to these simulated frequency space measurements. We take a 2D subsampled mask in the xyx-y-plane and sample fully in the $z$-direction as in [peper2020highly]. We consider different subsampling rates $R=\frac{N_{sampled}}{N_{total}}={8,16,32}$, with two different masks: the pseudo-spiral mask and the pseudo-random Gaussian mask, which is sampled according to a Gaussian probability distribution, as shown in Figure 2. For the pseudo-spiral mask, the points are placed evenly on a cartesian grid along a spiral with six turns and a final radius reaching the edge of the mask.
Additionally, we require measurements of the magnitude itself. The value of the background phase is treated as known instead.
@@ 3.1.3 Inverse problem setup
We consider two different initial guesses for these parameters, as in [garay2022]:
@@
The initial standard deviation for the reparameterized parameters was set to $0.5$, i.e. $P^{0}=0.5\mathbb{I}$, meaning that the prior models that there is $\approx 95\%$ probability that the target value will lie within the range of half/twice the initial guess. The standard deviation of the noise was estimated from the initial time step by computing the standard deviation of
$\boldsymbol{Y}^{0}-\boldsymbol{M}(t^{0})\odot\exp\left(i\phi_{back}(t^{0}))\right)\odot\boldsymbol{S}$
@@ 3.2.1 Flow phantom measurements
We used data from a flow experiment on a phantom of the carotid artery reported in [peper2020highly]. The phantom is made of a distensible silicon and is suspended in water. It takes the shape of a bifurcating tube, to simulate a carotid artery, and a backflow tube. A pump generates a pulsatile flow with a rate of roughy 60 bpm, simulating a cardiac cycle lasting 1s.
The 4D Flow MRI scan was performed with a 3T Ingenia scanner by Philips Healthcare. All three velocity directions were acquired, plus one acquisition with no encoding gradient to acquire the background phase. The scan parameters were set to TR = 8.9ms, TE = 4.5ms, FA = 8, vencvenc=150cm/s. The matrix size was [160, 160, 40] with a spatial resolution of [0.8mm, 0.8mm, 0.8mm] and a temporal resolution of 19 frames per cardiac cycle. The scan was accelerated with an acceleration factor of R=2. The scan used a 32-channel coil leading to 15 independent coil measurements, resulting in a total matrix size of $[160,160,40,4,15,19]$. Examples of the measurements, both the frequency space magnitude and different reconstructions, are shown in Figure 5.